Show that 𝑦 = 2𝑥 ln 𝑥 is a solution to the differential equation "x^2y''-xy'+y=0"



1
Expert's answer
2022-06-02T05:09:03-0400
"x^2y''-xy'+y=0"

"y=2x\ln x"

"y'=2\ln x+2"

"y''=2\/x"

"x^2y''-xy'+y=\\
=x^2(2/x)-2x\ln x-2x+2x\ln x=0"